OpenStudy (anonymous):

Fing the missing monomial -6x^10. ? =84x^23

ganeshie8 (ganeshie8):

say, the missing thing is A

ganeshie8 (ganeshie8):

-6x^10. A =84x^23

ganeshie8 (ganeshie8):

solve A

OpenStudy (anonymous):

Help me out with it plz

ganeshie8 (ganeshie8):

sure, you simply need to isolate A start by dividing -6x^10 both sides

ganeshie8 (ganeshie8):

-6x^10. A =84x^23 divide -6x^10 both sides 84x^23 A = -------- -6x^10

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

So whats the next steps

ganeshie8 (ganeshie8):

cancel watever u can

ganeshie8 (ganeshie8):

6 goes in 84 how many times ?

OpenStudy (anonymous):

6 divided by 84 is 14

ganeshie8 (ganeshie8):

good :) 84x^23 A = -------- -6x^10 14x^23 A = -------- -x^10

ganeshie8 (ganeshie8):

wat about x's ?

OpenStudy (anonymous):

Now I understand how you arrived then you subtract the exponent 10 and 23 to get 13

ganeshie8 (ganeshie8):

Yup !